OpenStudy (mayah3):

HELP PLEASE!!!!: 7 + f – 21 = –20

OpenStudy (oaktree):

Sorry, f-14=-20.

OpenStudy (oaktree):

So x=-6.

OpenStudy (oaktree):

I'm sure it's right.

OpenStudy (turingtest):

subtract 7 from both sides: f-21=-27 add 21 to both sides f-21+21=-27+21 f=-6